R is between Q and S, and S is between R and T. If QS=8, QT=14, and RT=11 find RS
Answer by ewatrrr(24785)


Hi
  Q  R  S  T 
QT- RT =  14 - 11 = QR = 3
 QS - 3 = 8 - 3 =  5 = RS

Wish You the Best in your Studies.